OpenStudy (steve816):

Warning: Extremely difficult. Please attempt this for me!!! Prove the identity.

OpenStudy (steve816):

\[1-8\sin^2\theta \cos^2\theta =\cos(4\theta)\]

OpenStudy (seascorpion1):

Use demoveries theorem.

OpenStudy (wolframwizard):

1-2sinθcosθ*2sinθcosθ*2 1-2sin(2θ)sin(2θ) 1-2sin^2(2θ) cos(2*2θ) cos(4θ)

OpenStudy (seascorpion1):

Great answer!!!

OpenStudy (steve816):

Wow @WolframWizard You're a genius!

OpenStudy (steve816):

May I ask what your thought process was? I just can't seem to understand how you can figure this out.

OpenStudy (wolframwizard):

I try to pull the expressions apart as much as possible so I can use identities to piece it back together into what I want.